OpenStudy (anonymous):

rains a and b are traveling in the same direction on parallel tracks train a is traveling 80 per hour and train b is traveling 84 miles per hour. train a passes a station at 8:15pm if train b passes the same station at 8:45 when will train b catch up with train a?

OpenStudy (anonymous):

Train a time1 = t hrs, Train b time2 = (t - 0.5) hrs. d2 = d1, speed2*time2=speed1*time1 84(t - 0.5) = 80t, 84t - 42 = 80t, 84t - 80t = 42, 4t = 42, t = 42/4 = 10.5 hrs, (t - 0.5) = 10.5 - 0.5 =10hrs to catchup.

OpenStudy (anonymous):

so the time would be 10:30 pm right?

OpenStudy (anonymous):

6:45 am

OpenStudy (anonymous):

hope it helps!